if f(x)=7x-5 and g(x)=x squared +9 what is fg(1)
Answer by stanbon(75887)

if f(x)=7x-5 and g(x)=x squared +9 what is fg(1)
------
fg(1) = f(1)*g(1) = (7-5)(1) = 2
==================================
If you meant fog(1) you would getf[g(1)] = f(1) = 7-5 = 2
========================
Cheers,
Stan H.